# Green Functions in Small Characteristic

### v1

@@ Abstract
In this note we solve the last open cases, which are for exceptional groups of type $E_{8}(q)$ where $q$ is a power of $2$, $3$ or $5$.
@@ 1 Introduction
Let $p$ be a prime, $q$ some power of $p$ and $\mathbb{F}_{q}$ the field with $q$ elements. Let $\mathbf{G}$ be a reductive algebraic group over an algebraic closure ${\bar{\mathbb{F}}_{p}}$ with a Frobenius endomorphism $F$ defining an $\mathbb{F}_{q}$-rational structure.
@@
In [Lus85] Lusztig defined another set of class functions on the unipotent elements using character sheaves, these are called generalized Green functions. A subset of them is also associated to $G(q)$-classes of $F$-invariant maximal tori, and Lusztig showed later [Lus90] that the two types of Green functions coincide.
@@
More precisely, this algorithm determines the Green functions as linear combination of certain functions which are supported on elements $C^{F}$ for a single unipotent class $C$ in $\mathbf{G}$. The values of these functions are clear up to a normalization by some scalar in $\mathbb{C}$ of absolute value $1$.
@@
So, with the results of this paper the Green functions are known in all cases. It only remains to consider the other generalized Green functions in groups of exceptional types.
Our method is a variant of the idea in [Gec20a] where the permutation character of the Borel subgroup in $G(q)$ and computations in a matrix representation where used. Here, we use more general parabolic subgroups and compute with the Steinberg presentation of the considered groups.
@@ 2 Notations
Let $\mathbf{T}\subset\mathbf{G}$ be an $F$-stable maximal torus of $\mathbf{G}$ that is contained in an $F$-stable Borel subgroup $\mathbf{B}\subseteq\mathbf{G}$. The group $\mathbf{G}$ is determined up to isomorphism by its root datum with respect to $\mathbf{T}$. The minimal unipotent subgroups $X_{r}$ normalized by $\mathbf{T}$ are called root subgroups. There is an isomorphism $\mathbf{k}^{+}\to X_{r}$, $a\mapsto x_{r}(a)$. The root $r:\mathbf{T}\to k^{\times}$ is an element of the (additive) character group $X(\mathbf{T})$ and decribes the conjugation: $tx_{r}(a)t^{-1}=x_{r}(r(t)a)$. We write $\Phi$ for the finite set of roots, then we have $\Phi=\Phi^{+}\dot{\cup}\Phi^{-}$ where a root $r$ lies in $\Phi^{+}$ when $X_{r}\subseteq\mathbf{B}$; these subsets are called positive roots and negative roots, respectively, and we have $\Phi^{-}=-\Phi^{+}$. The positive roots contain a unique subset $\Delta$ such that every positive root is a unique non-negative linear combination of the roots in $\Delta$; the elements of $\Delta$ are called simple roots. The height $\textrm{ht}(r)$ of a positive root $r\in\Phi^{+}$ is the sum of its coefficients when $r$ is written as linear combination of the roots in $\Delta$.
@@
The group $W=N_{\mathbf{G}}(\mathbf{T})/\mathbf{T}$ is called the Weyl group of $\mathbf{G}$, it is a Coxeter group with Coxeter generators $S:=\{s_{r}\mid\;r\in\Delta\}$, where $s_{r}$ is the unique non-trivial coset of $\mathbf{T}$ in $N_{\mathbf{G}}(T)\cap\langle X_{r},X_{-r}\rangle$. The length $l(w)$ of an element $w\in W$ is the smallest integer $k$ such that $w=s_{1}\cdots s_{k}$ with $s_{i}\in S$, the sequence $s_{1},\ldots,s_{k}$ is called a reduced word for $w$. An element $w\in W$ permutes the roots $r\in\Phi$ where $w(r)$ is defined by $wX_{r}w^{-1}=X_{w(r)}$. For $r\in\Delta$ we have $s_{r}(r)=-r$ and $s_{r}(\Phi^{+}\setminus\{r\})\subset\Phi^{+}$. So, for $w\in W$ and $r\in\Delta$ we have $l(ws_{r})<l(w)$ (and in that case $l(ws_{r})=l(w)-1$) if and only if $w(r)\in\Phi^{-}$.
@@
$(BsB)(BwB)\left\{\begin{array}[]{ll}=B(sw)B,&\textrm{ if }l(sw)>l(w)\\ \subseteq(BwB)\cup(B(sw)B),&else.\end{array}\right.$
@@ 3 Computing in the unipotent subgroup
We want to do certain computations with elements in $\mathbf{G}$ or $G(q)$ where the elements are represented in the strong form of the Bruhat decomposition. For this one first needs to fix the isomorphisms $k^{+}\to X_{r}$, $t\mapsto x_{r}(t)$ and the representatives $\dot{w}$ of Weyl group elements. And for the multiplication one needs to know commutator rules for the factors appearing in Bruhat expressions of the form $g=ut\dot{w}u^{\prime}$. These are provided by the relations in the Steinberg presentation of the group $\mathbf{G}$ or $G(q)$, see for example [Spr98, Ch.9].
@@ 3.2 Commutator formula
Furthermore, using the formulae in [Car72, 4.3.1,5.2.2] we can compute for any two different roots $r_{1},r_{2}\in\Phi^{+}$ and any $i,j\in\mathbb{Z}_{>0}$ such that $ir_{1}+jr_{2}\in\Phi^{+}$ an integer $C_{ijr_{1}r_{2}}$. These enable us to compute with elements in the group $U=\langle X_{r}\mid\;r\in\Phi^{+}\rangle$ using the following commutator formula [Car72, 5.2.3]:
@@
Let $r_{1},r_{2},\ldots r_{N}$ be the positive roots of $\mathbf{G}$ in any fixed order, we write $x_{i}(a):=x_{r_{i}}(a)$ for the corresponding root elements.
@@
    Any element of $U$ can be uniquely written in the form
@@
    Any product of root elements $x_{t_{1}}(b_{1})\cdots x_{t_{k}}(b_{k})$ for positive roots $r_{t_{1}},\ldots,r_{t_{k}}$ can be rewritten to the form in (a) (where some factors $x_{i}(0)=1$ may be omitted) by applying a finite number of the following steps to any pair of consecutive factors $x_{t_{i}}(b_{i})x_{t_{i+1}}(b_{i+1})$ with $t_{i+1}\leq t_{i}$:
@@
    Let $\prec$ be the partial ordering on the set of positive roots with $r_{i}\prec r_{j}$ if and only if $r_{j}-r_{i}$ is a non-zero non-negative linear combination of positive roots. Let $r_{i}$ be a positive root, and
@@
    Then we have $b_{i}=a_{i}+a$ and $b_{j}=a_{j}$ whenever $r_{i}\not\prec r_{j}$.
@@ 4 Green functions by the Lusztig-Shoji algorithm
Assume that the action of $F$ on $\Phi$ is trivial. Let $C$ be an F-stable unipotent class with representative $u\in C^{F}=C\cap G(q)$. Assume that $F$ acts trivially on $A(u)$ (so that we have $\tilde{\rho}=\rho$ in the discussion above). Let $\rho\in\textrm{Irr}(A(u))$ appear in the Springer correspondence and $\zeta_{u,\rho}$ be the associated scalar. Then we have for any positive integer $m$ that $u$ is also stable under $F^{m}$. Write $\zeta^{(m)}_{u,\rho}$ for the scalar associated to $(u,\rho)$ when $u$ is considered as element of $G^{F^{m}}=G(q^{m})$. Then we have $\zeta^{(m)}_{u,\rho}=(\zeta_{u,\rho})^{m}$.
What remains to find the Green functions $Q_{w}$ (or equivalently the $Q_{\tilde{\chi}}$ for $F$-stable irreducible $\chi$) as class functions is to choose representatives $u$ of unipotent classes and the characters $\tilde{\rho}$ and then to determine the scalars $\zeta_{u,\tilde{\rho}}$. As mentioned in Section 1 this task has been done in almost all cases. We describe a method which enabled us to handle the remaining cases of groups of type $E_{8}(q)$ in bad characteristic $2$, $3$ and $5$.
@@
From this we can always determine the sizes of the $G(q)$-classes in $C^{F}$ for each class: If for a class $C$ there are $k$ classes in $C^{F}$ and we write $Y_{C}$ for the $k\times k$ matrix of the values of the $\zeta_{u,\tilde{\rho}}Y_{u,\tilde{\rho}}$, and if $\Lambda_{C}$ is the corresponding diagonal block of $\Lambda$, then $Y_{C}\Lambda_{C}\bar{Y_{C}}^{t}$ is a diagonal matrix where the diagonal entries are the class lengths.
@@ 5.1 Permutation characters by Deligne-Lusztig characters
    The Green function $Q_{w}$ for the maximally split torus is the permutation character ${1_{B(q)}}^{G(q)}$. From its values we can find the sizes of the intersections of the unipotent classes in $G(q)$ with $B(q)$ or $U(q)$: For $u\in G(q)$ unipotent we have
@@
    We have $\sum\limits_{u\in G(q)\textrm{ {\scriptsize unipotent}}}Q_{w_{i}}(u)=\frac{G(q)}{T_{w_{i}}(q)}.$
We could also mention the orthogonality relations for Green functions, but in our applications they never provided useful information.
@@ 5.2 Permutation characters by counting fixed points
    $\{\dot{w}\prod_{r\in\Phi^{+},\,w(r)\in\Phi^{-}}x_{r}(a_{r})|\;w\textrm{ is $J$-reduced,}\;a_{r}\in k\}.$
@@
    $\{\dot{w}\prod_{r\in\Phi^{+},\,w(r)\in\Phi^{-}}x_{r}(a_{r})|\;w\textrm{ is $J$-reduced,}\;a_{r}\in\mathbb{F}_{q}\}.$
@@
Recall that $\mathbf{P}_{J}$ is generated by the $\mathbf{B}\dot{s}_{r}\mathbf{B}$ with $r\in J$. Let $g^{\prime}\in\mathbf{G}$ and $w^{\prime}\in W$ such that $g^{\prime}\in\mathbf{B}\dot{w}^{\prime}\mathbf{B}$. When $w^{\prime}$ is not $J$-reduced, there is $r\in J$ with $l(s_{r}w)<l(w)$ ($\Leftrightarrow w^{\prime-1}(r)\in\Phi^{-}$ by [Car72, 2.2.1]). We set $w=s_{r}w^{\prime}$ and have $(\mathbf{B}\dot{s}_{r}\mathbf{B})(\mathbf{B}\dot{w}\mathbf{B})=\mathbf{B}\dot{w}^{\prime}\mathbf{B}$. This shows that the coset $\mathbf{P}g^{\prime}=\mathbf{P}g$ for some $g\in\mathbf{B}\dot{w}\mathbf{B}$. Since the length of $w$ is smaller than the length of $w^{\prime}$ we will find after a finite number of applications of this step $g$ and $w$ such that $w$ is $J$-reduced. Since $\mathbf{B}\leq\mathbf{P}_{J}$ we see from the strong form of the Bruhat decomposition that the elements given in (a) contain representatives of all right cosets in $\mathbf{P}_{J}\setminus\mathbf{G}$.
@@ 5.3 Computing a character value
at once. For the computation we use independent indeterminates $y_{l+1},\ldots y_{N}$ over $\mathbb{F}_{q}$ (or even over $\mathbb{Z}$) instead of the $a_{i}$ in the expression above. We have to count for how many specializations $a_{i}\in\mathbb{F}_{q}$ of the $y_{i}$ the elements
@@
So, we want to count the tuples $(a_{l+1},\ldots,a_{N})\in\mathbb{F}_{q}^{N-l}$ such that specializing the $y_{i}$ by $a_{i}$ in the expressions
@@ 5.4 Counting solutions
  • Otherwise, we use a straight forward backtrack search: Specialize one variable to all possible values and solve for each value recursively the resulting system of equations in the remaining variables.
@@ 6 Application to $\mathbf{E_{8}(q)}$
As an application we want to determine the (ordinary) Green functions for the exceptional groups of type $E_{8}(q)$ in bad characteristic $p=2,3,5$. For good characteristic $p>5$ this problem was solved by Beynon and Spaltenstein in [BS85]. In Section 4 we have explained that we need to find appropriate class representatives of the unipotent classes in $\mathbf{G}$ and certain associated complex scalars $\zeta_{u,\bar{\rho}}$ of absolute value $1$. The following theorem summarizes the result.
@@
The exception is the class $C$ labeled $D_{8}(a_{3})$ when $q\equiv-1(3)$ (so $p\neq 3$). In this case there is a unique $G^{F}$-class of $u\in C^{F}$ such that $F$ acts trivially on $A(u)\cong S_{3}$. Then $\zeta_{u,\rho}=1$ except when $\rho=-1$ is the sign character of $S_{3}$ where $\zeta_{u,-1}=-1$.
@@
We can conclude from Mizuno’s results that all unipotent classes contain an $F$-stable element $u$ such that $F$ acts trivially on $A(u)$: Otherwise there would be a class $C$ such that $F$ acts on $A(u)$ as a non-inner automorphism for any $u\in C^{F}$. All groups occuring as $A(u)$ have the easy to check property that the number of $\phi$-conjugacy classes for any non-inner automorphism $\phi$ is smaller than the number of conjugacy classes. So, for some power $F^{m}$ of $F$ such that $F^{m}$ acts trivially on all $A(u)$ the group $G^{F^{m}}$ would have more unipotent classes than $G^{F}$. But Mizuno showed that the number of unipotent classes only depends on the characteristic $p$.
@@
In other cases certain possibilities for $\zeta_{u,\rho}$ do not lead to a diagonal matrix of class length as described in 4.2. For example for the class $D_{4}(a_{1})$, also with $A(u)\cong S_{3}$, we get that the two not yet known scalars can only be $1$, and this determines the class lengths.
An interesting case are classes $C$ where for $u\in C$ we have $A(u)$ of order $2$ and the two classes in $C^{F}$ have different length. There both choices in {±$\{\pm 1\}$ of the scalar for the non-trivial character of $A(u)$ lead to a diagonal matrix with both class lengths, but in different order. So, here we can just assume that the scalar is $\zeta_{u,-1}=1$ and we get the class length of $u$.
@@
Step (6). Two of the remaining four unknown scalars belong to the class $C=D_{8}(a_{3})$. For $u\in C$ we have $A(u)\cong S_{3}$, so that the $G(q)$-class of $u$ is fixed by the condition that $F$ acts trivially on $A(u)$. We still need to determine $\zeta_{u,-1}$ and $\zeta_{u,\rho}$ where we write $-1$ for the sign character and $\rho$ for the character of degree $2$ of $S_{3}$.
@@
Remarks: The computation of each value took about 5 seconds, there are $3628800$ $J$-reduced elements $w$, but only about $500$ of them do not fulfill the criterion in Section 5.5. Note that with our computation we can actually show that the elements $z_{52},z_{53},z_{54}$ in Mizuno’s list are indeed in the class $C=D_{8}(a_{3})$ because the same values do not occur on other classes. In fact, computing the value of $z_{44}$ in Mizuno’s list also yields the value $177889$ and this shows that this element is conjugate to $z_{53}$.
@@
As in step (6) we can use the parabolic subgroup for $J$ of type $D_{4}$. The four possibilities for the two unknown scalars always lead to the same four character values, but they differ in which of the four values appears twice. For $q=2$ the four values are $\{6785,2625,6401,2241\}$. Using the five representative $z_{i}$, $27\leq i\leq 31$, in Mizuno’s list and computing the character values for these elements we get the value $6785$ twice, namely for $z_{27}$ and $z_{31}$. This shows that also the scalars $\zeta_{u,\epsilon^{\prime}}=\zeta_{u,\epsilon^{\prime\prime}}=1$.

### v2

@@ Abstract
In this note we solve the last open cases, which are for exceptional groups $E_{8}(q)$ where $q$ is a power of $2$, $3$ or $5$.
@@ 1 Introduction
Let $p$ be a prime, $q$ some power of $p$ and $\mathbb{F}_{q}$ the field with $q$ elements. Let $\mathbf{G}$ be a connected reductive algebraic group over an algebraic closure ${\bar{\mathbb{F}}_{p}}$ with a Frobenius endomorphism $F$ defining an $\mathbb{F}_{q}$-rational structure.
@@
In [Lus85] Lusztig defined another set of class functions on the unipotent elements using character sheaves, these are called generalized Green functions. A subset of them is also associated to $G(q)$-classes of $F$-invariant maximal tori, and Lusztig [Lus90] (for large $q$) and Shoji [Sho95, 2.2] showed later that the two types of Green functions coincide.
@@
More precisely, this algorithm determines the Green functions as linear combinations of certain functions which are supported on elements $C^{F}$ for a single unipotent class $C$ in $\mathbf{G}$. The values of these functions are clear up to a normalization by some scalar in $\mathbb{C}$ of absolute value $1$.
@@
So, with the results of this paper the Green functions are known in all cases. It only remains to consider the other generalized Green functions in groups of exceptional types. This is done in [LS24], using the results of this article.
Our method is a variant of the idea in [Gec20a] where the permutation character of the Borel subgroup in $G(q)$ and computations in a matrix representation were used. Here, we use more general parabolic subgroups and compute with the Steinberg presentation of the considered groups.
Explicit tables of Green functions for exceptional groups are available from [Lüb24].
@@ 2 Notations
Let $\mathbf{T}\subseteq\mathbf{G}$ be an $F$-stable maximal torus of $\mathbf{G}$ that is contained in an $F$-stable Borel subgroup $\mathbf{B}\subseteq\mathbf{G}$. The group $\mathbf{G}$ is determined up to isomorphism by its root datum with respect to $\mathbf{T}$. The minimal unipotent subgroups $X_{r}$ normalized by $\mathbf{T}$ are called root subgroups. There is an isomorphism $\mathbf{k}^{+}\to X_{r}$, $a\mapsto x_{r}(a)$. The root $r:\mathbf{T}\to\mathbf{k}^{\times}$ is an element of the (additive) character group $X(\mathbf{T})$ and describes the conjugation: $tx_{r}(a)t^{-1}=x_{r}(r(t)a)$. We write $\Phi$ for the finite set of roots, then we have $\Phi=\Phi^{+}\dot{\cup}\Phi^{-}$ where a root $r$ lies in $\Phi^{+}$ when $X_{r}\subseteq\mathbf{B}$; these subsets are called positive roots and negative roots, respectively, and we have $\Phi^{-}=-\Phi^{+}$. The positive roots contain a unique subset $\Delta$ such that every positive root is a unique non-negative linear combination of the roots in $\Delta$; the elements of $\Delta$ are called simple roots. The height $\textrm{ht}(r)$ of a positive root $r\in\Phi^{+}$ is the sum of its coefficients when $r$ is written as linear combination of the roots in $\Delta$.
@@
The group $W=N_{\mathbf{G}}(\mathbf{T})/\mathbf{T}$ is called the Weyl group of $\mathbf{G}$, it is a Coxeter group with Coxeter generators $S:=\{s_{r}\mid\;r\in\Delta\}$, where $s_{r}$ is the unique non-trivial coset of $\mathbf{T}$ in $N_{\mathbf{G}}(\mathbf{T})\cap\langle X_{r},X_{-r}\rangle$. The length $l(w)$ of an element $w\in W$ is the smallest integer $k$ such that $w=s_{1}\cdots s_{k}$ with $s_{i}\in S$, the sequence $s_{1},\ldots,s_{k}$ is called a reduced word for $w$. An element $w\in W$ permutes the roots $r\in\Phi$ where $w(r)$ is defined by $wX_{r}w^{-1}=X_{w(r)}$. For $r\in\Delta$ we have $s_{r}(r)=-r$ and $s_{r}(\Phi^{+}\setminus\{r\})\subset\Phi^{+}$. So, for $w\in W$ and $r\in\Delta$ we have $l(ws_{r})<l(w)$ (and in that case $l(ws_{r})=l(w)-1$) if and only if $w(r)\in\Phi^{-}$.
@@
$(BsB)(BwB)\left\{\begin{array}[]{ll}=B(sw)B,&\textrm{if }l(sw)>l(w)\\ \subseteq(BwB)\cup(B(sw)B),&\textrm{otherwise.}\end{array}\right.$
@@ 3 Computing in the unipotent subgroup
We want to do certain computations with elements in $\mathbf{G}$ or $G(q)$ where the elements are represented in the strong form of the Bruhat decomposition. For this one first needs to fix the isomorphisms $\mathbf{k}^{+}\to X_{r}$, $t\mapsto x_{r}(t)$ and the representatives $\dot{w}$ of Weyl group elements. And for the multiplication one needs to know commutator rules for the factors appearing in Bruhat expressions of the form $g=ut\dot{w}u^{\prime}$. These are provided by the relations in the Steinberg presentation of the group $\mathbf{G}$ or $G(q)$, see for example [Spr98, Ch.9].
@@ 3.2 Commutator formula
Furthermore, using the formulae in [Car72, 4.3.1,5.2.2] we can compute for any two different roots $r_{1},r_{2}\in\Phi^{+}$ and any $i,j\in\mathbb{Z}_{>0}$ such that $ir_{1}+jr_{2}\in\Phi^{+}$ an integer $C_{ijr_{1}r_{2}}$. These enable us to compute with elements in the group $\mathbf{U}=\langle X_{r}\mid\;r\in\Phi^{+}\rangle$ using the following commutator formula [Car72, 5.2.3]:
@@
Let $r_{1},r_{2},\ldots,r_{N}$ be the positive roots of $\mathbf{G}$ in any fixed order, we write $x_{i}(a):=x_{r_{i}}(a)$ for the corresponding root elements.
@@
    Any element of $\mathbf{U}$ can be uniquely written in the form
@@
    Any product of root elements $x_{t_{1}}(b_{1})\cdots x_{t_{k}}(b_{k})$ where $r_{t_{1}},\ldots,r_{t_{k}}$ are positive roots can be rewritten to the form in (a) (where some factors $x_{i}(0)=1$ may be omitted) by applying a finite number of the following steps to any pair of consecutive factors $x_{t_{i}}(b_{i})x_{t_{i+1}}(b_{i+1})$ with $t_{i+1}\leq t_{i}$:
@@
    For positive roots $r_{i},r_{j}$ we write $r_{i}\prec r_{j}$ if and only if $r_{j}-r_{i}$ is a non-zero non-negative linear combination of positive roots.
    Let $r_{i}$ be a positive root, and
@@
    Then we have $b_{i}=a_{i}+a$ and $b_{j}=a_{j}$ whenever $j\neq i$ and $r_{i}\not\prec r_{j}$.
@@ 4 Green functions by the Lusztig-Shoji algorithm
Assume that the action of $F$ on $\Phi$ is trivial. Let $C$ be an F-stable unipotent class with representative $u\in C^{F}=C\cap G(q)$. Assume that $F$ acts trivially on $A(u)$ (so that we have $\tilde{\rho}=\rho$ in the discussion above). Let $\rho\in\textrm{Irr}(A(u))$ appear in the Springer correspondence and $\zeta_{u,\rho}$ be the associated scalar. Then we have for any positive integer $m$ that $u$ is also stable under $F^{m}$. Write $\zeta^{(m)}_{u,\rho}$ for the scalar associated to $(u,\rho)$ when $u$ is considered as element of $\mathbf{G}^{F^{m}}=G(q^{m})$. Then we have $\zeta^{(m)}_{u,\rho}=(\zeta_{u,\rho})^{m}$.
What remains to find the Green functions $Q_{w}$ (or equivalently the $Q_{\tilde{\chi}}$ for $F$-stable irreducible $\chi$) as class functions is to choose representatives $u$ of unipotent classes and the characters $\tilde{\rho}$ and then to determine the scalars $\zeta_{u,\tilde{\rho}}$. As mentioned in Section 1 this task has been done in almost all cases. We describe a method which enabled us to handle the remaining cases of groups $E_{8}(q)$ in bad characteristic $2$, $3$ and $5$.
@@
From this we can always determine the sizes of the $G(q)$-classes in $C^{F}$ for each class: If for a class $C$ there are $k$ classes in $C^{F}$ and we write $Y_{C}$ for the $k\times k$ matrix whose rows contain the values of the $\zeta_{u,\tilde{\rho}}Y_{u,\tilde{\rho}}$, and if $\Lambda_{C}$ is the corresponding diagonal block of $\Lambda$, then $Y_{C}^{-1}\Lambda_{C}\overline{Y_{C}}^{-t}$ is a diagonal matrix where the diagonal entries are the class lengths.
@@ 5.1 Permutation characters by Deligne-Lusztig characters
    The Green function $Q_{w}$ for the maximally split torus is the permutation character ${1_{B(q)}}^{G(q)}$ restricted to unipotent elements. From its values we can find the sizes of the intersections of the unipotent classes in $G(q)$ with $B(q)$ or $U(q)$: For $u\in G(q)$ unipotent we have
@@
    We have $\sum\limits_{u\in G(q)\textrm{ {\scriptsize unipotent}}}Q_{w_{i}}(u)=\frac{|G(q)|}{|T_{w_{i}}(q)|}.$
We could also mention the orthogonality relations for Green functions, but in our applications they never provided useful additional constraints.
@@ 5.2 Permutation characters by counting fixed points
    $\{\dot{w}\prod_{r\in\Phi^{+},\,w(r)\in\Phi^{-}}x_{r}(a_{r})\mid\;w\textrm{ is $J$-reduced,}\;a_{r}\in\mathbf{k}\}.$
@@
    $\{\dot{w}\prod_{r\in\Phi^{+},\,w(r)\in\Phi^{-}}x_{r}(a_{r})\mid\;w\textrm{ is $J$-reduced,}\;a_{r}\in\mathbb{F}_{q}\}.$
@@
Recall that $\mathbf{P}_{J}$ is generated by the $\mathbf{B}\dot{s}_{r}\mathbf{B}$ with $r\in J$. Let $g^{\prime}\in\mathbf{G}$ and $w^{\prime}\in W$ such that $g^{\prime}\in\mathbf{B}\dot{w}^{\prime}\mathbf{B}$. When $w^{\prime}$ is not $J$-reduced, there is $r\in J$ with $l(s_{r}w^{\prime})<l(w^{\prime})$ ($\Leftrightarrow w^{\prime-1}(r)\in\Phi^{-}$ by [Car72, 2.2.1]). We set $w=s_{r}w^{\prime}$ and have $(\mathbf{B}\dot{s}_{r}\mathbf{B})(\mathbf{B}\dot{w}\mathbf{B})=\mathbf{B}\dot{w}^{\prime}\mathbf{B}$. This shows that the coset $(\mathbf{P}_{J})g^{\prime}=(\mathbf{P}_{J})g$ for some $g\in\mathbf{B}\dot{w}\mathbf{B}$. Since the length of $w$ is smaller than the length of $w^{\prime}$ we will find after a finite number of applications of this step $g$ and $w$ such that $w$ is $J$-reduced. Since $\mathbf{B}\leq\mathbf{P}_{J}$ we see from the strong form of the Bruhat decomposition that the elements given in (a) contain representatives of all right cosets in $\mathbf{P}_{J}\setminus\mathbf{G}$.
@@ 5.3 Computing a character value
at once. For the computation we use independent indeterminates $y_{l+1},\ldots,y_{N}$ over $\mathbb{F}_{q}$ (or even over $\mathbb{Z}$) instead of the $a_{i}$ in the expression above. We have to count for how many specializations $a_{i}\in\mathbb{F}_{q}$ of the $y_{i}$ the elements
@@
So, we want to count the tuples $(a_{l+1},\ldots,a_{N})\in\mathbb{F}_{q}^{N-l}$ such that specializing the $y_{i}$ to $a_{i}$ in the expressions
@@ 5.4 Counting solutions
  • Otherwise, we use for fixed $q$ a straight forward backtrack search: Specialize one variable to all possible values and solve for each value recursively the resulting system of equations in the remaining variables.
@@ 6 Application to $\mathbf{E_{8}(q)}$
As an application we want to determine the (ordinary) Green functions for the exceptional groups $\mathbf{G}^{F}=G(q)=E_{8}(q)$ in bad characteristic $p=2,3,5$. For good characteristic $p>5$ this problem was solved by Beynon and Spaltenstein in [BS85]. In Section 4 we have explained that we need to find appropriate class representatives of the unipotent classes in $\mathbf{G}$ and certain associated complex scalars $\zeta_{u,\tilde{\rho}}$ of absolute value $1$. The following theorem summarizes the result.
@@
The exception is the class $C$ labeled $D_{8}(a_{3})$ when $q\equiv-1(3)$ (so $p\neq 3$). In this case there is a unique $\mathbf{G}^{F}$-class of $u\in C^{F}$ such that $F$ acts trivially on $A(u)\cong S_{3}$. Then $\zeta_{u,\rho}=1$ except when $\rho=-1$ is the sign character of $S_{3}$ where $\zeta_{u,-1}=-1$.
@@
We can conclude from Mizuno’s results that all unipotent classes contain an $F$-stable element $u$ such that $F$ acts trivially on $A(u)$: Otherwise there would be a class $C$ such that $F$ acts on $A(u)$ as a non-inner automorphism for any $u\in C^{F}$. All groups occurring as $A(u)$ have the easy to check property that the number of $\phi$-conjugacy classes for any non-inner automorphism $\phi$ is smaller than the number of conjugacy classes. So, for some power $F^{m}$ of $F$ such that $F^{m}$ acts trivially on all $A(u)$ the group $\mathbf{G}^{F^{m}}$ would have more unipotent classes than $\mathbf{G}^{F}$. But Mizuno showed that the number of unipotent classes only depends on the characteristic $p$.
@@
In other cases certain possibilities for $\zeta_{u,\rho}$ do not lead to a diagonal matrix of class lengths as described in 4.2. For example for the class $D_{4}(a_{1})$, also with $A(u)\cong S_{3}$, we get that the two not yet known scalars can only be $1$, and this determines the class lengths.
An interesting case are classes $C$ where for $u\in C$ we have $A(u)$ of order $2$ and the two classes in $C^{F}$ have different lengths. There both choices in {±$\{\pm 1\}$ of the scalar for the non-trivial character of $A(u)$ lead to a diagonal matrix with both class lengths, but in different order. So, here we can just assume that the scalar is $\zeta_{u,-1}=1$ and we get the class length of $u$.
@@
Step (6). Two of the remaining four unknown scalars belong to the class $C=D_{8}(a_{3})$. For $u\in C$ we have $A(u)\cong S_{3}$, so that the $G(q)$-class of $u$ is fixed by the condition that $F$ acts trivially on $A(u)$. We still need to determine $\zeta_{u,-1}$ and $\zeta_{u,\rho}$ where we write $-1$ for the sign character and $\rho$ for the irreducible character of degree $2$ of $S_{3}$.
@@
Remarks: The computation of each value took about 5 seconds, there are $3628800$ $J$-reduced elements $w$, but only about $500$ of them do not fulfill the criterion in Section 5.5. Note that with our computation we can actually show that the elements $z_{52},z_{53},z_{54}$ in Mizuno’s list are indeed in the class $C=D_{8}(a_{3})$ because the same values do not occur on other classes. In fact, we found an error in Mizuno’s list by computing the value of $z_{44}$. This also yields the value $177889$ and this shows that this element is conjugate to $z_{53}$.
@@
As in step (6) we can use the parabolic subgroup for $J$ of type $D_{4}$. The four possibilities for the two unknown scalars always lead to the same four character values, but they differ in which of the four values appears twice. For $q=2$ the four values are $\{6785,2625,6401,2241\}$. Using the five representatives $z_{i}$, $27\leq i\leq 31$, in Mizuno’s list and computing the character values for these elements we get the value $6785$ twice, namely for $z_{27}$ and $z_{31}$. This shows that also the scalars $\zeta_{u,\epsilon^{\prime}}=\zeta_{u,\epsilon^{\prime\prime}}=1$.
@@
Of course, using the techiques of this paper it is possible to check and recover known results for other types of groups of small rank (where the original argument involved much more elaborate computations).